Let the joint probability functions of X and Y be

f(x,y) = 2(x + y - 3xy2) , if 0 < x < 1, 0 < y < 1

0 , Otherwise

Find (i) Marginal probability density function of X and Y

(ii) Conditionally density functions


1
Expert's answer
2021-11-24T18:32:31-0500

(i)


"f_X(x)=\displaystyle\int_{-\infin}^{\infin}f(x, y)dy=\displaystyle\int_{0}^{1} 2(x + y - 3xy^2)dy"

"=2[xy+\dfrac{y^2}{2}-xy^3]\begin{matrix}
 1 \\
 0
\end{matrix}=2x+1-2x=1"

"f_X(x)=\begin{cases}
 1 & 0<x<1\\
 0 & otherwise
\end{cases}"



"f_Y(y)=\displaystyle\int_{-\infin}^{\infin}f(x, y)dx=\displaystyle\int_{0}^{1} 2(x + y - 3xy^2)dx"

"=2[\dfrac{x^2}{2}+xy-\dfrac{3x^2y^2}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}=1+2y-3y^2"

"f_Y(y)=\begin{cases}
 1+2y-3y^2 & 0<y<1\\
 0 & otherwise
\end{cases}"

(ii)

The conditional probability density function of "Y" given that "X = x" is


"f_{Y|X}=\dfrac{f(x, y)}{f_X(x)}=\dfrac{2(x + y - 3xy^2)}{1}"

"=2(x + y - 3xy^2),0 < x < 1, 0 < y < 1"

The conditional probability density function of "X" given that "Y=y" is


"f_{X|Y}=\dfrac{f(x, y)}{f_Y(y)}=\dfrac{2(x + y - 3xy^2)}{1+2y-3y^2} ,"

"0 < x < 1, 0 < y < 1"